OpenStudy (anonymous):

how can I find time in A=P(1+rt)?

OpenStudy (amistre64):

is that equation right?

OpenStudy (anonymous):

first, divide by P, so u get A/P=1+rt. then subtract 1, then divide by r. so u get (A/Pr)-1=t. i think that's it.

OpenStudy (amistre64):

if anything; \[\frac{A}{Pr}-\frac{1}{r}=t\]

OpenStudy (anonymous):

usually this formula is \[A=(1+r)^t\] just saying is all...

OpenStudy (amistre64):

i know right :) but you gotta take what your given

OpenStudy (anonymous):

^

OpenStudy (anonymous):

yes... master

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

(may i call you master?)

OpenStudy (amistre64):

.... call me whatever makes you feel best ;)

OpenStudy (anonymous):

in case it is \[A=(1+r)^t\] to solve for t you write \[t=\frac{\ln(A)}{\ln(1+r)}\]